Following is the information of a product of a firm:
Selling price per unit = $50
Variable expenses per unit = $20
Fixed expenses per month = $30,000
Desired operating income = $15,000

The breakeven point volume in units is ??
The volume necessary to earn the desired operating income is ??


{Please show calculations. I can't for the life of me figure this out! TIA!}
Answer #1

Contribution margin per unit

= Selling price per unit - Variable costs per unit

= 50 - 20

= 30 per unit.

A)

Breakeven point in volume = Fixed expenses/Contribution margin per unit

= 30,000/30

= 1,000 units

B)

Volume necessary = (Fixed cost + Desired income)/Contribution margin per unit

= (30,000+15,000)/30

= 1,500 units